# Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths

### v1

@@ 2.2 The mass-complete sample of HectoMAP quiescent galaxies
We explore the evolution in both average size and light concentration of massive quiescent galaxies, measured at fixed rest-frame wavelength in a mass-complete sample (Damjanov et al. 2023). We also examine the dependence of size and light concentration on galaxy stellar mass. A mass-complete sample is critically important for tracing evolutionary trends.
@@
To identify the quiescent population in the HectoMAP sample, Damjanov et al. (2023) select galaxies with $D_{n}4000>1.5$. The selection of the quiescent sample based on the $D_{n}4000$ indicator is equivalent to selection based on galaxy rest-frame colors (Damjanov et al. 2018). There are 51,274 quiescent galaxies in HectoMAP (seventh row of Table 1). The orange points in the central panel of Figure 1 show the $M_{\star}-z$ relation of HectoMAP quiescent galaxies. The orange histograms in the upper and right panels show the marginal distribution in $z$ and $M_{\star}$ for these galaxies.
@@
The HectoMAP quiescent sample is incomplete at $z<0.2$ (Sohn et al. 2021, 2023). As a consequence of the red-selection of HectoMAP the survey is missing a significant fraction of the low-mass quiescent galaxies that are bright enough to be included at low redshift. In Damjanov et al. (2023) we use the Smithsonian Hectospec Lensing Survey Field 2 sample (SHELS, Geller et al. 2014, 2016) to build a mass-complete quiescent sample from the HectoMAP parent sample. Although the SHELS F2 survey is slightly shallower than HectoMAP (with a limiting $r-$band magnitude $r_{\rm max}^{\rm SHELS}=20.75$), SHELS F2 is not color-selected. We translate the apparent magnitude limits of the HectoMAP and SHELS F2 samples into corresponding absolute magnitude limits as a function of redshift. Using the distribution of $M/L$ ratios for galaxies in narrow redshift bins, we then compute the redshift-dependent stellar mass limits for the two complete quiescent galaxy samples. The blue curve in Figure 1 shows the stellar mass limit for HectoMAP.
@@
We next use SHELS F2 for assessing the redshift interval where HectoMAP provides a complete mass-limited sample of quiescent galaxies. We compare mass distributions of SHELS F2 quiescent galaxies and all SHELS F2 galaxies with the HectoMAP color selection in $\Delta z=0.1$ redshift bins spanning $0.60.1<z<$ 0.6. For galaxies in the $0.60.2<z<$ 0.6 redshift range, the red color-selected SHELS F2 samples include 90%-100% of galaxies in the quiescent SHELS F2 samples. However in the $0.20.1<z<$ 0.2 redshift bin the color-selected sample is significantly smaller than the complete quiescent sample.
@@
We conclude that HectoMAP provides a complete mass-limited quiescent sample of 39,468 galaxies covering the redshift range $0.60.2<z<$ 0.6 (eighth and ninth entries in Table 1). We analyze this quiescent sample. In Figure 1, the quiescent galaxies (orange points) enclosed in the region outlined by the green solid lines constitute the HectoMAP mass-compete quiescent sample. The green histograms in the upper and right panels show the marginalized distribution along the corresponding axis. The median redshift and stellar mass of the sample (green lines in the upper and right panels) are $z=0.37_{-0.09}^{+0.09}$ and $\log(M_{\star}/M_{\odot})=10.87_{-0.24}^{+0.24}$, respectively. This large sample provides a powerful platform for examining the evolution of massive quiescent galaxy population.
@@ 3 Galaxy morphology from Subaru/HSC imaging
We characterize variation in the structural parameters of quiescent galaxies as a function of rest-frame wavelength. We also explore the dependence of these variations on redshift and stellar population age. We next describe the measurements of morphological parameters for the HectoMAP galaxies based on HSC-SSP multi-band imaging from Subaru/HSC. We summarize the HSC-SSP imaging of the HectoMAP field in Sect. 3.1; we describe the single-Sérsic model fitting procedure in Sect. 3.2; finally, and we report the results in Sect. 3.3.
@@ 3.1 HSC imaging
The HSC-SSP PDR3 (Aihara et al. 2022) provides $g\text{-,}$, $r\text{-,}$, $i\text{-,}$, $z\text{-,}$, and $-y\text{-}$band imaging of the HectoMAP region. The survey imaging covers the full HectoMAP field in the $g$-, $r$-, and $z$-bands, and 41 deg2 ($\sim$80of the survey) in the $i$- and $y$-band.
The image quality of HSC-SSP is excellent. The median seeing (i.e., the full width at half maximum (FWHM) of the PSF) is $0\aas@@fstack{\prime\prime}83$, $0\aas@@fstack{\prime\prime}74$, $0\aas@@fstack{\prime\prime}55$, $0\aas@@fstack{\prime\prime}62$, and $0\aas@@fstack{\prime\prime}69$ for the $g\text{-,}$, $r\text{-,}$, $i\text{-,}$, $z\text{-,}$, and $-y\text{-}$ band, respectively. In addition to the wavelength dependence of atmospheric seeing, different visiting times for different bands and changes in observing conditions also contribute to variations in image quality. The seeing is remarkably stable across the survey area: in each band, the local seeing is generally within $\sim 0\aas@@fstack{\prime\prime}2$ of the median value [^1] For HSC-SSP quality assurance plots, see https://hsc-release.mtk.nao.ac.jp/doc/index.php/quality-assurance-plots__pdr3/..
The hscPipe system (Bosch et al. 2018) is the standard pipeline for HSC-SSP (Miyazaki et al. 2018). The pipeline facilitates reduction of individual chips, mosaicking, and image stacking. The HSC-SSP DR3 offers two sets of coadded images based on two different approaches to sky subtraction: local and global (see Figure 8 of Aihara et al. 2022). We select coadds with local subtraction to mitigate the contribution of the extended wings of bright stars that may contaminate background galaxies. Only bright sources lose their wings with local sky subtraction. Sources with an extent $\lesssim 30$ arcsec ($\sim 100-200$ kpc at $0.60.2<z<$ 0.6) remain unaffected (Aihara et al. 2022). The sample we analyze does not contain galaxies that approach this size limit (see Sect. 3.3).
@@ 3.2 Single-Sérsic model fitting pipeline
where $I_{0}$ is the central intensity (surface brightness), $R_{e}$ is the half-light radius along the galaxy major axis, and $n$ is the Sérsic index, which describes the concentration of the model profile. The coefficient $b(n)$, a function of $n$, ensures that $R_{e}$ encloses half of the total galaxy light. The coefficient $b(n)$ is accurately approximated as (Ciotti & Bertin 1999)
$b(n)=2n-\frac{1}{3}+\frac{4}{405n}+\frac{46}{25515n^{2}}+\frac{131}{1148175n^{3}}.$ (2)
Extraction of Sérsic model parameters involves three steps. First, we run the SExtractor software (Bertin & Arnouts 1996) to obtain a catalog that includes detection of point-like sources. Then, we run the PSFex software (Bertin 2011) on this catalog to construct a set of spatially-varying PSFs. The PSF model at the position of an extended source is a linear combination of basic vectors that best fits the observed point-source profiles in the vicinity of that position. Finally, we run the SourceXtractor++ software (Bertin et al. 2020; Kümmel et al. 2020) [^2] Comprehensive SourceXtractor++ documentation is in https://sextractor.readthedocs.io/en/latest/index.html. to fit the galaxy surface brightness profile with a PSF-convolved Sérsic model (Eq. (1)) in 2D. Following this procedure we build the catalog of morphological parameters for the HectoMAP galaxy sample.
SourceXtractor++ , the successor to the original SExtractor package, fits models to sources by minimizing a least-squares loss function with respect to a parameter vector describing the shape of the 2D Sérsic model. SourceXtractor++ includes a flexible and Python-based model configuration, the possibility of defining complex priors and dependencies for model parameters, and code optimization for fast processing.
Euclid Collaboration et al. (2023) analyze the performance of the SourceXtractor++ model fitting by using realistic simulations of galaxies in Euclid’s visible band $I_{E}$. This broad band covers the range $\approx(5400-9000)\,\AA$ (Laureijs et al. 2011), overlapping with the $r$, $i$, and $z$ bands of HSC. For single-Sérsic model fitting to realistic synthetic profiles, the software provides reliable structural measurements with a bias $\lesssim 15\%$ and a scatter $\lesssim 30\%$ for apparent magnitude $I_{E}\sim 23$. Model uncertainties are typically underestimated by a factor $\sim 2$. Euclid Collaboration et al. (2023) test SourceXtractor++ along with other source-extraction and fitting software packages developed for large-area imaging surveys and conclude that SourceXtractor++ is the best approach.
The SourceXtractor++ 2D Sérsic model fitting provides best-fit values for seven free parameters: the 2D cartesian coordinates of the galaxy center, the ellipse rotation angle $\theta$, the ellipse minor-axis to major-axis ratio $b/a$, along with $I_{0}$, $R_{e}$, and $n$. Fits to the spatial model parameters are in pixel units which are then converted to arcseconds.
We apply a standard initialization for the seven fitting parameters:
@@
  • $b/a=0.5$, allowing exponential variation within the range $(0.1,1)$ ;
  • isophotal guess for the central surface brightness $I_{0}$, allowing exponential variation within ±$\pm 10^{3}$ times the initial proxy (from the built-in function get_flux_parameter);
  • isophotal half-light radius for $R_{e}$, for exponentially-spaced bins within (0.1,10)×(0.1,10$\times$( 0.1 , 10 ) × the isophotal value interval;
  • $n=1$, allowing linear variation in the range $(0,10)$.
We also consider a mild nongaussian prior for the Sérsic parameter $n$ to minimize the chance of obtaining (large) final values of $n\sim 9-10$. We require that the random variable $N=\exp{\left(n-n_{0}\right)}$, with $n_{0}=8$, is normally distributed with zero mean and standard deviation equal to 1. This prior is constant (uninformative) for $n\lesssim 7$, and decreases for larger $n$: it assigns a starting probability of $\sim 1/2$ to $n\sim 8$ and a probability of $\sim 0$ to $n\gtrsim 9$.
@@ 3.3 Sérsic Fitting Results
For each of the five HSC photometric bands, we measure Sérsic morphological parameters for all of the galaxies in the HectoMAP sample with HSC-SSP PDR3 imaging. We obtain best-fit models for all HectoMAP targets in the region with images available in all five bands ($\sim 78,000$ galaxies, or 82% of the parent HectoMAP sample; Sect. 3.1, second row in Table 1).
SourceXtractor++ identifies HectoMAP galaxies in HCS-SSP by requiring that the fitted model centroid lie within a radius of five pixels from the input centroid listed in the HectoMAP datatbase. The angular distance between the model and input centroid is $<50\%$ of the best ($i$-band) PSF FWHM for $>99\%$ of measured galaxies (Sect. 3.1). After identifying all of the targets in the HectoMAP catalog, SourceXtractor++ proceeds to fit a 2D Sérsic model to each galaxy surface brightness profile (Sect. 3.2). Table 2 lists the resulting best-fit parameters $R_{e}$, $b/a$, and $n$, along with $z$ and $M_{\star}$.
@@
Notes. This table is available in its entirety in a machine-readable form in the online journal. A portion is shown here for guidance regarding its form and content.
@@
The median reduced $\chi^{2}$ of the fits is in the range $(1.3-6.7)$, with the 25th, 50th, and 75th percentile of the reduced $\chi^{2}$ of $(1.1,1.5,3.0)$, $(1.4,2.8,9.1)$, $(2.4,6.7,26.0)$, $(1.3,2.5,8.9)$, and $(1.1,1.3,2.6)$ for the HSC $g\text{-,}$, $r\text{-,}$, $i\text{-,}$, $z\text{-,}$, and $-y\text{-}$band, respectively. The average goodness-of-fit magnitude $\chi^{2}\sim 1$ confirms small deviations between the best-fit single-Sérsic models and the light profiles of quiescent HectoMAP galaxies. In Appendix A we further compare the $i-$band fits with independent estimates of morphological parameters derived by fitting the HSC $i-$band images of the HectoMAP quiescent sample using SExtractor Damjanov et al. (2023). The agreement is excellent between the two sets of measurements.
@@
Figure 2: Half-light radius $R_{e}$ and Sérsic index $n$ in five HSC bands for 26,912 quiescent galaxies. These galaxies have morphological measurements in all five bands and $R_{e}$ exceeds half od the seeing FWHM.. Central panel: distributions of $-r\text{-}$ band in the $R_{e}-n$ plane. Galaxies (points) are segregated into two classes of $D_{n}4000$: $1.61.5<D_{n}4000<$ 1.6 (dark blue) and $D_{n}4000>1.6$ (green). Black contours show the probability density of HectoMAP galaxies in the $R_{e}-n$ plane derived from a Gaussian kernel estimator (KDE). The largest to the smallest contour enclose regions where the probability density exceeds the 10th to the 90th percentiles of the distribution. The upper and right panels show the marginal distributions in $R_{e}$ and $n$, respectively, color-coded by the HSC band. The equivalently color-coded dashed lines indicate the medians of these distributions. To improve readability, the central panel displays a randomly selected $30\%$ of the total sample. Histograms and isocurves are based on the total samples.
@@
The large panel in Figure 2 shows the distribution of measurements in the $R_{e}-n$ plane for the $-r\text{-}$band. Results for the other bands are qualitatively similar. For spheroidal galaxies with $n\gtrsim 2$ ($\sim 85-95\%$ of the sample) the expected positive correlation between galaxy size and concentration is evident (e.g., Caon et al. 1993; Trujillo et al. 2001; Schombert 2013; Vika et al. 2015). Less centrally concentrated galaxies (with $n\sim 1$) have larger radii along the semi-major axis ($R_{e}$) on average. There is an anticorrelation between size and concentration.
@@
Nearly all spheroidal galaxies ($n>2$) have old stellar populations ($D_{n}4000>1.6$, green points in Figure 2). Quiescent galaxies with younger stellar age ($1.61.5<D_{n}4000<$ 1.6, blue points in the figure) populate the small-$n$ and large-$R_{e}$ (bottom right) region of the $R_{e}-n$ plane. These galaxies, newcomers to the quiescent population (Sects. 2.2 and 4.2), drive the anticorrelation between galaxy size and Sérsic index.
@@
The histograms in the upper and right panels of Figure 2 show the distributions of $R_{e}$ and $n$ for the final sample of quiescent galaxies, color-coded by the HSC band. The distributions show similar shapes across all five bands. Sizes generally decrease as the wavelength increases: the median $\log(R_{e}/$ kpc)))) is $\sim 0.83,0.76,0.73,0.72,$ and $0.75$ in the $g\text{-,}$, $r\text{-,}$, $i\text{-,}$, $z\text{-,}$, and $-y\text{-}$band, respectively. In contrast, Sérsic indices increase with increasing wavelength: the median $n$ is $\sim 3.6,3.7,3.8,4.0,$ and $4.3$ from $g\text{-,}$, to $-y\text{-}$band.
The upper (right) panel in Figure 2 shows a small secondary peak in the size (Sérsic index) distribution for the $g$-band, at $\log(R_{e}/$ kpc)1.21.6$\sim$ 1.2-1.6) ∼ 1.2 - 1.6 ($n\sim 8-9$). Only $\sim 100$ galaxies are responsible for this feature. We investigate the fit residuals for these galaxies; there is no significant difference with respect to the rest of the population. The $g$-band has the largest median seeing, $\sim 0.83$ arcsec. This lower image quality could cause overestimation of $R_{e}$ and $n$ for the most centrally concentrated galaxies. Because this issue affects only $\lesssim 0.5\%$ of the sample, the overall results are unaffected.
4 Multi-wavelength Sérsic profiles

### v2

@@ 2.2 The mass-complete sample of HectoMAP quiescent galaxies
We explored the evolution in both average size and light concentration of massive quiescent galaxies, measured at fixed rest-frame wavelength in a mass-complete sample (Damjanov et al. 2023). We also examined the dependence of size and light concentration on galaxy stellar mass. A mass-complete sample is critically important for tracing evolutionary trends.
To identify the quiescent population in the HectoMAP sample, Damjanov et al. (2023) selected galaxies with $D_{n}4000>1.5$. This index is a robust indicator of quiescence for several reasons. First, the narrow wavelength range of the $D_{n}4000$ measurement (Sect. 2.1) minimizes its dependence on dust. Second, the selection of quiescent samples based on $D_{n}4000$ is not overly sensitive to changes in stellar metallicity (e.g., Kauffmann et al. 2003; Le Borgne et al. 2006). Using the 30-band photometry and the MMT/Hectospec spectroscopy of galaxies in the COSMOS field, Damjanov et al. (2018) showed that the selection of quiescent samples at $z\lesssim 0.5$ based on $D_{n}4000$ is essentially equivalent to selection based on the rest-frame UVJ colors derived from photometric measurements over a very broad UV-infrared wavelength range. These advantages make the cut in $D_{n}4000$ a preferred method for selecting quiescent galaxies in spectro-photometric studies at $z<1$ (e.g. Vergani et al. 2008; Woods et al. 2010; Geller et al. 2014; Damjanov et al. 2018, 2019, 2023; Figueira et al. 2024).
@@
There are 51,274 quiescent galaxies in HectoMAP (seventh row of Table 1). The orange points in the central panel of Fig. 1 show the $M_{\star}-z$ relation of HectoMAP quiescent galaxies. The orange histograms in the upper and right panels show the marginal distribution in $z$ and $M_{\star}$ for these galaxies.
@@
The HectoMAP quiescent sample is incomplete at $z<0.2$ (Sohn et al. 2021, 2023). As a consequence of the red selection of HectoMAP, the survey is missing a significant fraction of the low-mass quiescent galaxies that are bright enough to be included at low redshift. In Damjanov et al. (2023) we use the Smithsonian Hectospec Lensing Survey Field 2 sample (SHELS, Geller et al. 2014, 2016) to build a mass-complete quiescent sample from the HectoMAP parent sample. Although the SHELS F2 survey is slightly shallower than HectoMAP (with a limiting $r$-band magnitude $r_{\rm max}^{\rm SHELS}=20.75$), SHELS F2 is not color-selected. We translate the apparent magnitude limits of the HectoMAP and SHELS F2 samples into corresponding absolute magnitude limits as a function of redshift. Using the distribution of $M/L$ ratios for galaxies in narrow redshift bins, we then compute the redshift-dependent stellar mass limits for the two complete quiescent galaxy samples. The blue curve in Fig. 1 shows the stellar mass limit for HectoMAP.
@@
Next, we used SHELS F2 for assessing the redshift interval where HectoMAP provides a complete mass-limited sample of quiescent galaxies. We compared the mass distributions of SHELS F2 quiescent galaxies and all SHELS F2 galaxies with the HectoMAP color selection in $\Delta z=0.1$ redshift bins spanning $0.1<z<0.6$. For galaxies in the $0.2<z<0.6$ redshift range, the red color-selected SHELS F2 samples include 90%-100% of galaxies in the quiescent SHELS F2 samples. However in the $0.1<z<0.2$ redshift bin the color-selected sample is significantly smaller than the complete quiescent sample.
@@
After removing galaxies at $0.4<z<0.42$, where a prominent night sky line significantly impacts $D_{n}4000$ measurements (Damjanov et al. 2025), we concluded that HectoMAP provides a complete mass-limited quiescent sample of 39,468 galaxies covering the redshift range $0.2<z<0.6$ (eighth and ninth entries in Table 1). Next, we analyzed this quiescent sample. In Fig. 1, the quiescent galaxies (orange points) enclosed in the region outlined by the green solid lines constitute the HectoMAP mass-compete quiescent sample. The green histograms in the upper and right panels show the marginalized distribution along the corresponding axis. The median redshift and stellar mass of the sample (green lines in the upper and right panels) are $z=0.37_{-0.09}^{+0.09}$ and $\log(M_{\star}/M_{\odot})=10.87_{-0.24}^{+0.24}$, respectively. This large sample provides a powerful platform for examining the evolution of massive quiescent galaxy population.
@@ 3 Galaxy morphology from Subaru/HSC imaging
We characterize variation in the structural parameters of quiescent galaxies as a function of rest-frame wavelength. We also explore the dependence of these variations on redshift and stellar population age. We then describe the measurements of morphological parameters for the HectoMAP galaxies based on HSC-SSP multiband imaging from Subaru/HSC. We summarize the HSC-SSP imaging of the HectoMAP field in Sect. 3.1; we describe the single-Sérsic model fitting procedure in Sect. 3.2; finally, we report the results in Sect. 3.3.
@@ 3.1 HSC imaging
The HSC-SSP PDR3 (Aihara et al. 2022) provides $g\text{-,}$ $r\text{-,}$ $i\text{-,}$ $z\text{-,}$ and $y\text{-}$band imaging of the HectoMAP region. The survey imaging covers the full HectoMAP field in the $g$, $r$, and zzbands, and 41 deg2 ($\sim$80% of the survey) in the $i$ and $y$ bands.
The image quality of HSC-SSP is excellent. The median seeing (i.e., the full width at half maximum (FWHM) of the PSF) is $0\aas@@fstack{\prime\prime}83$, $0\aas@@fstack{\prime\prime}74$, $0\aas@@fstack{\prime\prime}55$, $0\aas@@fstack{\prime\prime}62$, and $0\aas@@fstack{\prime\prime}69$ for the $g$, $r$, $i$, $z$, and $y$ band, respectively. In addition to the wavelength dependence of atmospheric seeing, different visiting times for different bands and changes in observing conditions also contribute to variations in image quality. The seeing is remarkably stable across the survey area: in each band, the local seeing is generally within $\sim 0\aas@@fstack{\prime\prime}2$ of the median value [^1] For HSC-SSP quality assurance plots, see https://hsc-release.mtk.nao.ac.jp/doc/index.php/quality-assurance-plots__pdr3/..
The hscPipe system (Bosch et al. 2018) is the standard pipeline for HSC-SSP (Miyazaki et al. 2018). The pipeline facilitates reduction of individual chips, mosaicking, and image stacking. The HSC-SSP DR3 offers two sets of coadded images based on two different approaches to sky subtraction: local and global (see Fig. 8 of Aihara et al. 2022). We select coadds with local subtraction to mitigate the contribution of the extended wings of bright stars that may contaminate background galaxies. Only bright sources lose their wings with local sky subtraction. Sources with an extent $\lesssim 30$ arcsec ($\sim 100-200$ kpc at $0.2<z<0.6$) remain unaffected (Aihara et al. 2022). The sample we analyze does not contain galaxies that approach this size limit (see Sect. 3.3).
@@ 3.2 Single-Sérsic model fitting pipeline
where $I_{0}$ is the central intensity (surface brightness), $R_{e}$ is the half-light radius along the galaxy major axis, and $n$ is the Sérsic index, which describes the concentration of the model profile. The coefficient $b(n)$, a function of $n$, ensures that $R_{e}$ encloses half of the total galaxy light (Ciotti & Bertin 1999).
Extraction of Sérsic model parameters involves three steps. First, we run the SExtractor software (Bertin & Arnouts 1996) to obtain a catalog that includes detection of point-like sources. Then, we run the PSFex software (Bertin 2011) on this catalog to construct a set of spatially varying PSFs. The PSF model at the position of an extended source is a linear combination of basic vectors that best fits the observed point-source profiles in the vicinity of that position. Finally, we run the SourceXtractor++ software (Bertin et al. 2020; Kümmel et al. 2020) [^2] Comprehensive SourceXtractor++ documentation is in https://sextractor.readthedocs.io/en/latest/index.html. to fit the galaxy surface brightness profile with a PSF-convolved Sérsic model (Eq. (1)) in 2D. Following this procedure, we built the catalog of morphological parameters for the HectoMAP galaxy sample.
SourceXtractor++ , the successor to the original SExtractor package, fits models to sources by minimizing a least-squares loss function with respect to a parameter vector describing the shape of the 2D Sérsic model. SourceXtractor++ includes a flexible and Python-based model configuration, the possibility of defining complex priors and dependences for model parameters, and code optimization for fast processing.
Euclid Collaboration et al. (2023) analyzed the performance of the SourceXtractor++ model fitting by using realistic simulations of galaxies in Euclid’s visible band $I_{E}$. This broadband range covers $\approx(5400-9000)\penalty 10000\$Å (Laureijs et al. 2011), overlapping with the $r$, $i$, and $z$ bands of HSC. For single-Sérsic model fitting to realistic synthetic profiles, the software provides reliable structural measurements with a bias $\lesssim 15\%$ and a scatter $\lesssim 30\%$ for apparent magnitude $I_{E}\sim 23$. Model uncertainties are typically underestimated by a factor $\sim 2$. Euclid Collaboration et al. (2023) test SourceXtractor++ along with other source-extraction and fitting software packages developed for large-area imaging surveys and conclude that SourceXtractor++ is the best approach.
The SourceXtractor++ 2D Sérsic model fitting provides best-fit values for seven free parameters: the 2D cartesian coordinates of the galaxy center, the ellipse rotation angle $\theta$, the ellipse minor-axis to major-axis ratio $b/a$, along with $I_{0}$, $R_{e}$, and $n$. Fits to the spatial model parameters are given in pixel units, which are then converted to arcseconds. We applied a standard initialization for the seven fitting parameters:
@@
  • $b/a=0.5$, allowing for exponential variation within the range $(0.1,1)$;
  • isophotal guess for the central surface brightness $I_{0}$, allowing for exponential variation within $\pm 10^{3}$ times the initial proxy (from the built-in function get_flux_parameter);
  • isophotal half-light radius for $R_{e}$, for exponentially spaced bins within $(0.1,10)\times$ the isophotal value interval;
  • $n=1$, allowing for linear variation in the range $(0,10)$.
We also considered a mild non-Gaussian prior for the Sérsic parameter, $n$, to minimize the chance of obtaining (large) final values of $n\sim 9-10$. We require the random variable $N=\exp{\left(n-n_{0}\right)}$, with $n_{0}=8$, to be normally distributed with zero mean and standard deviation equal to 1. This prior is constant (uninformative) for $n\lesssim 7$, and decreases for larger $n$: it assigns a starting probability of $\sim 1/2$ to $n\sim 8$ and a probability of $\sim 0$ to $n\gtrsim 9$.
@@ 3.3 Sérsic fitting results
For each of the five HSC photometric bands, we measured the Sérsic morphological parameters for all of the galaxies in the HectoMAP sample with HSC-SSP PDR3 imaging. We obtained best-fit models for all HectoMAP targets in the region with images available in all five bands ($\sim 78,000$ galaxies, or 82% of the parent HectoMAP sample; Sect. 3.1, second row in Table 1).
SourceXtractor++ identifies HectoMAP galaxies in HCS-SSP by requiring that the fitted model centroid lie within a radius of five pixels from the input centroid listed in the HectoMAP datatbase. The angular distance between the model and input centroid is $50\%$ of the best ($i$ band) PSF FWHM for $99\%$ of measured galaxies (Sect. 3.1). After identifying all of the targets in the HectoMAP catalog, SourceXtractor++ proceeds to fit a 2D Sérsic model to each galaxy surface brightness profile (Sect. 3.2). Table 2 lists the resulting best-fit parameters $R_{e}$, $b/a$, and $n$, along with $z$ and $M_{\star}$.
@@
Notes. This table is available in its entirety at the CDS. A portion is shown here for guidance regarding its form and content.
@@
The median reduced $\chi^{2}$ of the fits is in the range (1.36.7)(1.3-6.7), with the 25th, 50th, and 75th percentile of the reduced $\chi^{2}$ of $(1.1,1.5,3.0)$, $(1.4,2.8,9.1)$, $(2.4,6.7,26.0)$, $(1.3,2.5,8.9)$, and $(1.1,1.3,2.6)$ for the HSC $g$, $r$, $i$, $z$, and $y$ band, respectively. The average goodness-of-fit magnitude $\chi^{2}\sim 1$ confirms small deviations between the best-fit single-Sérsic models and the light profiles of quiescent HectoMAP galaxies. In Appendix A we further compare the $i$-band fits with independent estimates of morphological parameters derived by fitting the HSC $i$-band images of the HectoMAP quiescent sample using SExtractor Damjanov et al. (2023). The agreement is excellent between the two sets of measurements.
@@
Figure 2: Half-light radius, $R_{e}$, and Sérsic index, $n$, in five HSC bands for 26,912 quiescent galaxies. These galaxies have morphological measurements in all five bands and $R_{e}$ exceeds half of the seeing FWHM. The central panel shows the distributions of $r\text{-}$ band in the $R_{e}-n$ plane. Galaxies (points) are segregated into two classes of $D_{n}4000$: $1.5<D_{n}4000<1.6$ (dark blue) and $D_{n}4000>1.6$ (green). Black contours show the probability density of HectoMAP galaxies in the $R_{e}-n$ plane derived from a Gaussian kernel estimator (KDE). The largest to the smallest contour enclose regions where the probability density exceeds the 10th to the 90th percentiles of the distribution. The upper and right panels show the marginal distributions in $R_{e}$ and $n$, respectively, color-coded by the HSC band. The equivalently color-coded dashed lines indicate the medians of these distributions. To improve readability, the central panel displays a randomly selected $30\%$ of the total sample. Histograms and isocurves are based on the total samples.
@@
The large panel in Fig. 2 shows the distribution of measurements in the $R_{e}-n$ plane for the $r\text{-}$band. Results for the other bands are qualitatively similar. For spheroidal galaxies with $n\gtrsim 2$ ($\sim 85-95\%$ of the sample) the expected positive correlation between galaxy size and concentration is evident (e.g., Caon et al. 1993; Trujillo et al. 2001; Schombert 2013; Vika et al. 2015). Less centrally concentrated galaxies (with $n\sim 1$) have larger radii along the semi-major axis ($R_{e}$) on average. There is an anticorrelation between size and concentration.
@@
Nearly all spheroidal galaxies ($n>2$) have old stellar populations ($D_{n}4000>1.6$, green points in Fig. 2). Quiescent galaxies with younger stellar age ($1.5<D_{n}4000<1.6$, blue points in the figure) populate the small-$n$ and large-$R_{e}$ (bottom right) region of the $R_{e}-n$ plane. These galaxies, newcomers to the quiescent population (Sects. 2.2 and 4.2), drive the anticorrelation between galaxy size and Sérsic index.
[…]
@@
The histograms in the upper and right panels of Fig. 2 show the distributions of $R_{e}$ and $n$ for the final sample of quiescent galaxies, color-coded by the HSC band. The distributions show similar shapes across all five bands. Sizes generally decrease as the wavelength increases: the median $\log(R_{e}/$ kpc)) is $\sim 0.83,0.76,0.73,0.72,$ and $0.75$ in the $g$, $r$, $i$, $z$, and $y$ band, respectively. In contrast, Sérsic indices increase with increasing wavelength: the median $n$ is $\sim 3.6,3.7,3.8,4.0,$ and $4.3$ from the $g\text{}$ to the $y\text{ }$band.
A small fraction ($\sim 9.3\%$) of galaxies in our quiescent sample have $n>8$ in all bands (the secondary peaks in the right panel of Fig. 2). On average, galaxies with $n>8$ are $\sim 0.26$ mag fainter than galaxies with $n<8$. The fit residuals for $n>8$ galaxies show no significant difference with respect to the rest of the population. Moreover, the average reduced $\chi^{2}$ for galaxies with $n>8$ is within the interquartile range of the $\chi^{2}$ distribution for the subsample with $n<8$. We perform the analysis from Sects. 4 and 5 with and without high-$n$ galaxies and obtain indistinguishable results. To provide a comprehensive analysis, we report the results using the complete mass-limited sample.
[…]
4 Multiwavelength Sérsic profiles